OpenStudy (anonymous):

Tran and Thuy are packing supplies for a camping trip. They decide to pour their box of pancake mix into two separate bags.If Tran pours 2 1/3 cups of mix from the box, which contains 5 1/4 cups, how many cups are left in the box?

OpenStudy (anonymous):

box has 5 1/4 cups tran pours 2 1/3 cups what's left is: 5 1/4 - (2 1/3) = 5+1/4 - 2 - 1/3 = 5-2 + 1/4 - 1/3 = 3 + 3/12 - 4/12 = 3 - 1/12 = 2 + 12/12 - 12/12 = 2 + 11/12 = 2 11/12

OpenStudy (anonymous):

awsome :) <3

OpenStudy (anonymous):

I actually have a typo in one of the lines above. Can you identify it ?

OpenStudy (anonymous):

um....let me see.... i think its 5+1/4 - 2 - 1/3..